Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)

μ = 15.0; σ = 3.9

P(10 ≤ x ≤ 26) =
Answer #1

Here,

X ~ Normal(15, 3.92)

Now,

P(10< x < 26) = P(X < 26 - P(X < 10)

X - 1526 - 15 → P(10 < X < 26) = P6 15 ) 26 – 15 3.9 3, 95 3.9 3.9

—P(10 < x < 26) = P(Z < 2.82051) - P(Z < -1.28205)

= P(10 < X < 26) = (2.82051)-0(-1.28205)

= P(10 < X < 26) = 0.997603 - 0.0999123

P(10 < X < 26) = 0.897691

P(10 < X < 26) 0.8977